OpenStudy (anonymous):

An object moves along the path x^2 - y^2 = 5. At the point (-3, 2), dx/dt=-4. Find dy/dt at that same point.

OpenStudy (caozeyuan):

from the equation you can differentiate x on both sides and get dy/dx=f(x,y), which is actually a number is x and y are known

OpenStudy (caozeyuan):

now, dy/dx=dy/dt*dt/dx, since dt/dx=1/(dx/dt), there is only one unkown which is dy/dt, solve it and you are done

OpenStudy (anonymous):

Umm I'm still confused..

OpenStudy (caozeyuan):

\[2x-2yy'=0\] correct?

OpenStudy (anonymous):

Yes :)

OpenStudy (xapproachesinfinity):

yes that's correct

OpenStudy (xapproachesinfinity):

now solve for dy/dt

OpenStudy (xapproachesinfinity):

but you missed dx/dt

OpenStudy (caozeyuan):

\[y'=\frac{ x }{ y }\]

OpenStudy (caozeyuan):

if x=-3,y=2, y'=-3/2

OpenStudy (xapproachesinfinity):

x and y are all depend on t

OpenStudy (caozeyuan):

so dy/dt*dt/dx=-3/2

OpenStudy (caozeyuan):

since dx/dt=-4,dt/dx=1/-4

OpenStudy (xapproachesinfinity):

\(x^2-y^2=5\) \(2xx'-2yy'=0\)

OpenStudy (caozeyuan):

dy/dt*-1/4=-3/2, dy/dt=6

OpenStudy (caozeyuan):

we have diff notation, my y' is dy/dx, not dy/dt

OpenStudy (xapproachesinfinity):

solving for y' \(y'=xx'/y\)

OpenStudy (xapproachesinfinity):

now substitute

OpenStudy (anonymous):

Ohhhhh! Now I understand how to do dy/dt! :)

OpenStudy (caozeyuan):

and we arrive at same answer, but diff method

OpenStudy (xapproachesinfinity):

why are looking dy/dx? the question is asked to find dy/dt

OpenStudy (anonymous):

That makes so much more sense thank you guys! :)

OpenStudy (xapproachesinfinity):

eh i just realized @Breee615 is the poster haha thought it was you

OpenStudy (caozeyuan):

because I can use chain rule

OpenStudy (xapproachesinfinity):

yes you can! so you diff with respect to x first the implicit equation is that what you did?

OpenStudy (phi):

using x' is ambiguous, but works as long as you know it means dx/dt But being more explicit , and using implicit differentiation with respect to t: \[ \frac{d}{dt} \left(x^2 - y^2 = 5\right) \\2x \frac{dx}{dt} - 2y \frac{dy}{dt}= 0 \] now fill in the known values, and solve for dy/dt

OpenStudy (xapproachesinfinity):

yeah true, i wanted to write that way but kind lazy to use too much latex hahah

OpenStudy (caozeyuan):

yes, that exactly what I did, then I find out the dy/dx= some number by pluging in x and y, finally the dx/dt comes in during chain rule

OpenStudy (xapproachesinfinity):

but certainly i agree with you

OpenStudy (anonymous):

So wait my answer is 6 right?? If it isn't then I must not understand how to do it :o

OpenStudy (caozeyuan):

Yes it is

OpenStudy (xapproachesinfinity):

oh makes sense :)

OpenStudy (xapproachesinfinity):

you should have used dy/dx instead of y' hahah

OpenStudy (anonymous):

Okay good, thank you :)

OpenStudy (xapproachesinfinity):

that's chain rule still

OpenStudy (xapproachesinfinity):

so how did you get -12

OpenStudy (xapproachesinfinity):

the answer 6 for sure haha

OpenStudy (xapproachesinfinity):

the first step is correct old man :) what the next step

OpenStudy (xapproachesinfinity):

oh that's the parametric thingy

OpenStudy (xapproachesinfinity):

oh great job! you just did what @coazeyuan did just you nailed it down more

OpenStudy (xapproachesinfinity):

you used the parametric derivative dy/dx

OpenStudy (xapproachesinfinity):

you can use several variable calc but the poster should understand it though

OpenStudy (anonymous):

Do you guys mind helping me with another problem quick? :)

OpenStudy (xapproachesinfinity):

post in a diff post

OpenStudy (anonymous):

Okay! :)

OpenStudy (xapproachesinfinity):

and tag